write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0.
Solution:9, only solution

Answer by stanbon(75887)

write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2+bx+c=0.
Solution:9, only solution
============================
The quadratic is y = (x-9)^2
y = x^2 - 18x + 81
==============================
Cheers,
Stan H.